For interference in thin films, obtain the condition of constructive and destructive
interference between reflected light waves.
1
Expert's answer
2019-05-30T09:07:59-0400

If the wave fall on the film at the angle Θ, then the condition when the interference is constructive:

"2dn \cos{\Theta}=m\lambda, \space m=0,1,2.."

where n- refraction index of the film, d- the height of the film.

The interference is destructive if:


"2dn \cos{\Theta}=(m+1/2)\lambda, \space m=0,1,2.."